A gas is filled with oxygen at a temperature of 25°c to a pressure of 2atm. If the temperature is increased to 50°c and the pressure is increased to 5atm, by what factor does the density of the oxygen increase by.

1
Expert's answer
2022-02-23T08:31:25-0500

"\rho=\frac{pM}{RT},"

"\frac{\rho_2}{\rho_1}=\frac 52\cdot \frac{298}{323}=2.3."